OpenStudy (anonymous):

How do you convert 270 degrees to radians?? Need help steps please

OpenStudy (ipwnbunnies):

To convert degrees to radians, multiply by pi/180. For example: \[90 degrees*\frac{\pi}{180} = \frac{\pi}{2} radians\]

OpenStudy (anonymous):

so the anwser is 3pie/2 radians

OpenStudy (ipwnbunnies):

pi*

OpenStudy (unklerhaukus):

this works because pi radians equals 180° \[\pi[\text{rad}]=180°\\~\\ \frac{\pi[\text{rad}]}{180°}=1\] \[270°=270°\times1=270°\times\frac{\pi[\text{rad}]}{180°}=\frac{270}{180}\pi[\text{rad}]\]

OpenStudy (anonymous):

@UnkleRhaukus, you got the answer right but you know you could still have broken down 270/180 to 3/2.

OpenStudy (anonymous):

in Maths, every answer needs to be reduced to its lowest form

OpenStudy (unklerhaukus):

yes , but its not my job to answer the original question , i only try to help the op with the most difficult part, and leave the easy parts for them to solve by themselves, OS is not a 'do-someone-elses-homework-for-them' site, it is a 'help-some-with-their-homework' site

OpenStudy (unklerhaukus):

posting direct answers without working won't teach the op anything